# A First Look at Package-to-Group Mechanism: An Empirical Study of the Linux Distributions

### v1

@@ VII-B2 Key information in package descriptions
We extract the most relevant keywords that we believe may distinguish packages with P2G from those without it. Figure 11 shows the results using two clouds. We select the top frequent keywords based on TF-IDF and compare packages with P2G to those without. As shown, the keywords of packages with P2G emphasize general utility and development resources, such as “files”, “fonts”, “library”, “applications” and “development”. These packages focus more on infrastructure-related tasks and provide various supporting files for diverse development environments. On the other hand, packages without P2G highlight more technical aspects, with keywords like “debug”, “documentation”, “crate” and “bindings”. This suggests a focus on debugging tools, development documentation, and specific programming language frameworks. These packages are more developer-oriented, catering primarily to those involved in debugging and toolchain development.
@@ VIII-A Implications
Guidance for Developers and Maintainers: By identifying the categories of packages that commonly adopt the P2G(e.g., graphical desktop and network), developers and maintainers can better predict and decide which packages should adopt the P2G in future versions. This proactive categorization provides practical guidelines for package grouping, facilitating better decision-making in package management.
Support for OSS Project Evolution: The identification of evolutionary patterns in the P2G mechanism, such as group splitting and merging, allows developers to anticipate changes and better manage updates in large-scale OSS projects. Recognizing that packages no longer adopting the P2G mechanism are likely to remain within distributions highlights the resilience of OSS ecosystems in maintaining critical components.
Enhanced Quality Control through GVALUE: The development and validation of the GVALUE metric provide a concrete method to assess group quality within OSS projects. Identifying and addressing issues such as poor group descriptions and insufficient package numbers will lead to better-quality package groupings and enhance the long-term sustainability of the P2G mechanism. This insight can be applied beyond Linux distributions to other OSS projects that rely on third-party packages.
@@ VIII-B Threat to validity
Construct Validity concerns the relationship between the treatment and the outcome. The threat comes from the rationality of the questions asked. We are interested in assessing the package-to-groups mechanism in Linux distributions. To achieve this goal we focus on its evolving trend, change pattern, quality measurement, and developers’ tendency. We believe that these questions have a high potential to provide unique insights and value for practitioners and researchers.
Internal Validity addresses potential threats to the way the study was conducted. The first threat relates to the identification of packages and groups. This was done manually, which may have resulted in some omissions, but ensured the overall quality of the dataset. The second threat concerns the selection of increased and decreased groups when answering RQ2. While we analyzed a subset of samples, the margin of error remains within acceptable limits. The third threat relates to the topic extracted through topic modeling. We acknowledge that the choice of these topics is to some extent subjective. For example, different researchers might reach different results. To mitigate this threat, we invited two Linux practitioners to conduct the topic selection. After this initial proposal, we compared our topics and revised the results.
External Validity considers the generalizability of our findings. This study focused exclusively on popular open-source Linux distributions. Although the P2G mechanism is important across all OSS projects, our research revealed that it is rarely implemented in less popular projects. Furthermore, developers in such projects often lack the time to apply P2G, which may explain the limited application of P2G in these cases.
@@ IX Conclusion
In this paper, we conduct a preliminary study on the package-to-group mechanism from its application trend, evolutionary patterns, group quality, and developers’ tendencies. We find that although the P2G mechanism has become more prevalent, especially in popular distributions, it still represents a small proportion of the total packages. We also identify six major change patterns of P2G and find packages that no longer use the P2G tend to remain in Linux distributions. Additionally, we propose the GValue metric to evaluate the quality of groups, finding that many groups still exhibit poor quality. We also summarize types of packages that may be more inclined to adopt the P2G by Linux developers to support future updates. Our work investigates this mechanism from multiple aspects, which can provide significant insights for OSS developers and maintainers.

### v2

@@ VII-B2 Key information in package descriptions
Figure 11 shows the keywords that can distinguish packages with P2G from those without it, presented in two clouds. We can observe that the keywords of packages with P2G emphasize general utility and development resources, e.g., “files”, “fonts”, and “applications”. It suggest that packages with P2G focus more on infrastructure-related tasks and provide various development environments. The packages without P2G highlight more technical aspects, with keywords like “debug” and “documentation”. This suggests they focus more on debugging tools and development documentation.
@@ VIII-A Implications
Guidance for Developers and Maintainers: By identifying the categories of packages that commonly adopt P2G (e.g., graphical desktop and network), developers and maintainers can better predict and decide which packages should adopt P2G in future versions. This categorization can provide practical guidelines for them.
Support for OSS Project Evolution: Summarizing the evolution patterns of the P2G mechanism, such as splitting and merging groups, enables developers to predict potential changes and manage updates more effectively. The conclusion that packages that no longer adopt P2G are likely to remain within distributions underscores the resilience and long-term stability of OSS projects.
Enhanced Quality Control: The GVALUE provides a concrete method to assess the quality of groups within OSS projects. Identified issues (e.g., inappropriate descriptions and insufficient numbers) will help developers and maintainers to improve the quality of groups. The GVALE and insights can contribute to control and optimize groups and improve long-term project success.
@@ VIII-B Threat to validity
Construct Validity concerns the relationship between the treatment and the outcome. The threat comes from the rationality of the research questions we asked. We are interested in the package-to-groups mechanism in Linux distributions. To achieve this goal we focus on its application trend, evolution pattern, quality assessment, and group tendency. We believe that these questions have a high potential to provide unique insights and value for practitioners and researchers.
Internal Validity addresses potential threats to the way the study was conducted. The first threat relates to the identification of packages and groups. This was done manually, which may have resulted in some omissions, but ensured the overall quality of the dataset. The second threat concerns the manual summary of group change patterns when answering RQ2. To address this, we employ an experienced team and set a two-round labeling process. The third threat relates to the topic extracted through topic modeling. We acknowledge that the choice of these topics is to some subjective. To mitigate this threat, we invite two Linux practitioners to conduct the topic selection. After this initial proposal, we compare our topics and revised the results.
External Validity considers the generalizability of our findings. This study primarily focuses on popular open-source Linux distributions. Although the P2G mechanism is important across all OSS projects, our research reveals that it is rarely implemented in less popular projects (Section IV). Furthermore, developers working on these less active projects often face time constraints, which may hinder the adoption of P2G. This also explains the limited application in these cases.
@@ IX Conclusion
In this paper, we conduct a preliminary study on the package-to-group mechanism from its application trend, evolution patterns, group quality, and group tendency. We find that although the mechanism has become more prevalent, especially in popular distributions, it still represents a small proportion of the total packages. We also summarize six major change patterns of P2G and find packages that no longer use P2G tend to remain in Linux distributions. Additionally, we propose the GValue metric to evaluate the quality of groups, finding that many groups still exhibit poor quality. We also summarize types of packages that may be more inclined to adopt P2G by developers to support future updates. Our work investigates this mechanism from multiple aspects, which can provide significant insights for OSS projects developers and maintainers.
X ACKNOWLEDGMENT
This research was supported by the National Natural Science Foundation of China (62192731, 62192730, 62162051).